OpenStudy (anonymous):

6 + 7 * log2 x = 14 with 2 being the base of the log

OpenStudy (anonymous):

log2x = 8/7 x = 2 ^ 8/7 = 2.208

OpenStudy (anonymous):

2x=2^(8/7) x=1.1

OpenStudy (radar):

I think the 2 is not 2x but the base for the log to use.

OpenStudy (radar):

7* log (base 2) x = 14-6=8 8=2^3 7*3=21

OpenStudy (radar):

a swag i admit